# SoliReward: Mitigating Susceptibility to Reward Hacking and Annotation Noise in Video Generation Reward Models

### v1

@@ 4.2.2 Bradley-Terry with Win-Tie
Task Approach RM ACC ID OOD Phy & Deform Linear (Ln) [zhao2025swift] 74.69 78.66 ‘Yes’ token logits [wu2025rewarddance] 75.43 78.46 Special token + Ln [liu2025improving] 75.91 73.61 HPQA (Ours) 78.48 80.08 TA Linear (Ln) [zhao2025swift] 72.41 31.92 ‘Yes’ token logits [wu2025rewarddance] 71.56 31.37 Special token + Ln [liu2025improving] 76.25 58.38 HPQA (Ours) 79.02 60.25
@@ 4.3 Post-Training
We conduct post-training to further validate our reward model on physical plausibility and subject deformity. HunyuanVideo [kong2024hunyuanvideo] is selected as the video generation backbone. We apply online reinforcement learning post-training with DanceGRPO [xue2025dancegrpo], guided by our reward model. This process enhances video quality in these two aspects. To quantify this, we employ the VideoAlign Motion Quality (MQ) [liu2025improving] and VBench2 Human Fidelity [zheng2025vbench] evaluation metrics. As shown in  Tab. 4, the results confirm the efficacy of our reward model during this post-training phase.
@@ 4.4 Ablation Study
Model Architecture. We benchmark our HPQA against three common alternative architectures: 1) Extracting the last token embedding from the final hidden state and projecting it via a linear layer to derive a score; 2) VideoAlign [liu2025improving], which introduces a special token placeholder in the model input and uses its corresponding output embedding, processed by a linear layer, as the score; and 3) RewardDance [wu2025rewarddance], which instructs the reward model to output ‘yes’ or ‘no’ and utilizes the predicted probability of the ‘yes’ token as the reward. As  Tab. 5 indicates, our HPQA method outperforms the other architectures in terms of accuracy, validating its efficacy.
@@
Binary Cross Entropy. RewardDance [wu2025rewarddance] augments the BT loss with a cross-entropy penalty during point-wise RM training. However, it does not thoroughly explore this component’s impact on RM accuracy and subsequent post-training performance. Motivated by this gap, we empirically investigate the effect of this binary cross-entropy (BCE) penalty, treating the RM’s output as logits for a binary pass/fail task. The results reveal a critical insight. RM accuracy is a deceptive metric when decoupled from the reward margin. As shown in Tab. 6, adding the BCE penalty (‘BT-WT + BCE’) yields comparable accuracy to our ‘BT-WT’ baseline. However, this BCE-penalized RM leads to inferior post-training performance.
@@ 7.3.1 Physical Plausibility and Subject Deformity
Table 10: Comparison of different approaches on human preference benchmarks on VBench2 [zheng2025vbench]. The metrics are: Anatomy (Human Anatomy), Clothes (Human Clothes), Identity (Human Identity), and Fidelity (Human Fidelity).
@@ 7.3.2 Semantic Alignment
Table 13: Comparison of post-training with different reward models on semantic alignment. Evaluations are conducted on VBench [huang2024vbench].
@@ 8.2 Data Distribution
We collected annotations for $250\mathrm{k}$ in-house videos and a $50\mathrm{k}$-sample out-of-distribution (OOD) test set generated by other SOTA models (including Wan2.1 [wan2025wan], Wan2.2 [wan2025wan], Veo 3 [google2025veo] and Seedance 1.0 [gao2025seedance]). These samples originate from $20\mathrm{k}$ unique prompts. To ensure diversity, we balanced these prompts across multiple attributes including subject, motion, style, and camera control. In Fig. 7, we demonstrate the quality distribution of our in-house dataset. To construct the BT-WT dataset, we randomly selected 350kk win-lose and 150kk win-tie pairs based on the cross-prompt pairing strategy.
@@ 9 BT, BTT and BT-WT
As discussed in Sec. 3.3, VideoAlign [liu2025improving] retains all tie samples, using a loss function that models A-wins, B-wins, or ties (which includes both win-tie and lose-tie pairs). We contend this approach is ill-suited for our annotation scenario. Specifically, if two negative samples are labeled independently (e.g., for deformity), we cannot assume their degree of severity is equivalent. Erroneously treating them as a ‘tie’ (a lose-tie pair) would introduce label noise, diminishing the RM’s discriminative capacity and ultimately degrading policy performance.

### v2

@@ 4.2.2 Bradley-Terry with Win-Tie
Task Approach RM ACC ID OOD Phy & Deform Linear (Ln) [38] 74.69 78.66 ‘Yes’ token logits [32] 75.43 78.46 Special token + Ln [17] 75.91 73.61 HPQA (Ours) 78.48 80.08 TA Linear (Ln) [38] 72.41 31.92 ‘Yes’ token logits [32] 71.56 31.37 Special token + Ln [17] 76.25 58.38 HPQA (Ours) 79.02 60.25
@@ 4.3 Post-Training
We conduct post-training to further validate our reward model on physical plausibility and subject deformity. HunyuanVideo [12] is selected as the video generation backbone. We apply online reinforcement learning post-training with DanceGRPO [36], guided by our reward model. This process enhances video quality in these two aspects. To quantify this, we employ the VideoAlign Motion Quality (MQ) [17] and VBench2 Human Fidelity [39] evaluation metrics. As shown in  Tab. 4, the results confirm the efficacy of our reward model during this post-training phase.
@@ 4.4 Ablation Study
Model Architecture. We benchmark our HPQA against three common alternative architectures: 1) Extracting the last token embedding from the final hidden state and projecting it via a linear layer to derive a score; 2) VideoAlign [17], which introduces a special token placeholder in the model input and uses its corresponding output embedding, processed by a linear layer, as the score; and 3) RewardDance [32], which instructs the reward model to output ‘yes’ or ‘no’ and utilizes the predicted probability of the ‘yes’ token as the reward. As  Tab. 5 indicates, our HPQA method outperforms the other architectures in terms of accuracy, validating its efficacy.
@@
Binary Cross Entropy. RewardDance [32] augments the BT loss with a cross-entropy penalty during point-wise RM training. However, it does not thoroughly explore this component’s impact on RM accuracy and subsequent post-training performance. Motivated by this gap, we empirically investigate the effect of this binary cross-entropy (BCE) penalty, treating the RM’s output as logits for a binary pass/fail task. The results reveal a critical insight. RM accuracy is a deceptive metric when decoupled from the reward margin. As shown in Tab. 6, adding the BCE penalty (‘BT-WT + BCE’) yields comparable accuracy to our ‘BT-WT’ baseline. However, this BCE-penalized RM leads to inferior post-training performance.
@@ 7.3.1 Physical Plausibility and Subject Deformity
Table 10: Comparison of different approaches on human preference benchmarks on VBench2 [39]. The metrics are: Anatomy (Human Anatomy), Clothes (Human Clothes), Identity (Human Identity), and Fidelity (Human Fidelity).
@@ 7.3.2 Semantic Alignment
Table 13: Comparison of post-training with different reward models on semantic alignment. Evaluations are conducted on VBench [10].
@@ 8.2 Data Distribution
We collected annotations for $250\mathrm{k}$ in-house videos and a $50\mathrm{k}$-sample out-of-distribution (OOD) test set generated by other SOTA models (including Wan2.1 [29], Wan2.2 [29], Veo 3 [8] and Seedance 1.0 [7]). These samples originate from $20\mathrm{k}$ unique prompts. To ensure diversity, we balanced these prompts across multiple attributes including subject, motion, style, and camera control. In Fig. 7, we demonstrate the quality distribution of our in-house dataset. To construct the BT-WT dataset, we randomly selected 350kk win-lose and 150kk win-tie pairs based on the cross-prompt pairing strategy.
@@ 9 BT, BTT and BT-WT
As discussed in Sec. 3.3, VideoAlign [17] retains all tie samples, using a loss function that models A-wins, B-wins, or ties (which includes both win-tie and lose-tie pairs). We contend this approach is ill-suited for our annotation scenario. Specifically, if two negative samples are labeled independently (e.g., for deformity), we cannot assume their degree of severity is equivalent. Erroneously treating them as a ‘tie’ (a lose-tie pair) would introduce label noise, diminishing the RM’s discriminative capacity and ultimately degrading policy performance.